# Data Contamination Issues in Brain-to-Text Decoding

### v2

@@ 3.3 Our Method
To eliminate data contamination from both cognitive signal leakage and text stimuli leakage, we split the dataset by $\left\langle S_{i},T_{j}\right\rangle$ pairs as shown in (f) of Figure 2. Since EEG and fMRI are different in the composition of dataset, we treat them separately and propose two dataset splitting methods. As to EEG dataset where $F_{ij}$ and $T_{j}$ form a sample, we consider a bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$ where $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{T_{j}\}_{j=1}^{M}$. $\mathcal{E}$ is the edge between node in $\mathcal{U}$ and node in $\mathcal{V}$, indicating $\left\langle S_{i},T_{j}\right\rangle$ pair in the dataset. $N$ is the total number of subjects and $M$ is the total number of unique text stimuli. We assert $M>N$, so $e=(u,v)\in\mathcal{E}$ exists for every $v\in\mathcal{V}$, as each text stimuli is listened or read by at least one subject. As shown in step 2 of Figure 3, first we pick one edge for each node $v\in\mathcal{V}$ and build a new bipartite graph $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$. Then we split graph $\mathcal{G}_{2}$ by subject $\mathcal{U}$ with the given splitting ratio and form three disjoint graphs $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$. In step 4, some edges satisfying zero data contamination condition are not included in the graph. We add these edges to corresponding graphs, extending each graph $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ to its maximally scalable state and finishing the dataset splitting process.
@@
$F_{ij}=concat(f_{ij},f_{i,j+1},\ldots,f_{i,j+L-1})$ and $T_{j}=concat(s_{j},s_{j+1},\ldots,s_{j+L-1})$ form a sample pair in fMRI dataset. If we follow the same process as EEG, text stimuli leakage will occur in the overlapping part of two samples, when one sample is assigned to training set and the other is assigned to validation or test set. We propose a simple solution that achieves the balance between abandoning as little data as possible and ensuring zero data contamination. Instead of $\left\langle S_{i},T_{j}\right\rangle$ pair, we consider $\left\langle S_{i},M_{k}\right\rangle$ pair and apply the above-mentioned algorithm. More details and pseudo-code are available in Appendix B.
@@
Table 2: Generation quality of UniCoRN model for fMRI under different training settings. Here $*{}^{*}$ indicates the encoder of UniCoRN is randomly initialized instead of pre-trained through signal reconstruction task.
4 Experiments and Analysis
We test state-of-the-art brain-to-text decoding models on two popular cognitive datasets. Comprehensive experiments are conducted to prove the existence of the following phenomena: (1) Cognitive signals and text stimuli in test set leak into training set in all current dataset splitting methods. (2) The model’s generalization ability, particularly that of the auto-regressive decoder, has been overestimated due to data contamination. Because the number of tasks in EEG dataset is too small and method (e) makes no difference to EEG as method (d), we only consider method (a), (c), (d).
4.1 Datasets
@@
4.2 Implementation
We follow the same settings of UniCoRN (Xi et al., 2023) and EEG2Text (Wang and Ji, 2022), except the datasets are split to the ratio of 8:1:1 for fair comparison. All experiments are conducted on NVIDIA A100-SXM4-40GB GPUs. More details are shown in Appendix A.
[…]
$N=1$ $N=2$ $N=3$ $N=4$ $F$ $P$ $R$
@@ 4.2 Implementation
4.3 Data Contamination
We have analyzed two kinds of data contamination, cognitive signal leakage and text stimuli leakage in Methodology section. In this part, we will quantify data contamination situation through experiments.
To better illustrate the extent of data contamination across different dataset splitting methods, we design two novel evaluation metrics named Cognitive Signal Leakage Rate (CSLR) and Text Stimuli Leakage Rate (TSLR) for detecting cognitive signal leakage and text stimuli leakage. Note that the situation for validation set is similar as test set, we only consider test set in experiments. CSLR indicates the average percentage of each subject’s cognitive signals in test set appearing in training set, which could be formulated as
@@ 4.3 Data Contamination
where $N$ stands for the total number of subjects in test set. ||$\cdot$|| ⋅ | stands for the cardinality of a set. Function $\min(\cdot,\cdot)$ is applied to make sure for each subject the data leakage rate is less than 1.
The definition of TSLR is somewhat different for EEG signal and fMRI signal. As to EEG signal where cognitive signals are sampled continuously, it’s easy to match certain sentence stimuli with corresponding signals. Its TSLR is similar to CSLR, which indicates the average percentage of certain text in test set appearing in training set. TSLR for EEG data can be calculated through
@@
The fMRI signal is sampled discretely with a deterministic interval TR, making it hard to acquire signals corresponding to sentences. Previous methods instead concatenated continuous fMRI frames of certain length with their corresponding sentence segments as training samples. As a result, we consider the average percentage of the same sentence segments in test set appearing in training set as TSLR of fMRI signal. It can be formulated as
@@
We test current dataset splitting methods and our method on fMRI dataset Narratives and EEG dataset ZuCo. Considering the influence of randomness in splitting, we select four seeds for experiments. The results are shown in Table 1 and are consistent with theoretical analysis. For fMRI, current methods apart from method (a) suffer from cognitive signal leakage, while method (a) has serious text stimuli leakage. Method (b) gets no text stimuli leakage but has slight cognitive signal leakage. The situation for EEG is similar to that of fMRI. Apart from our proposed method (f), there is no way to achieve zero cognitive signal leakage and text stimuli leakage at the same time.
@@
4.4 Overestimated Performance
@@
Cognitive signal leakage and text stimuli leakage affect models from encoder side and decoder side.
4.4.1 Encoder
@@
As shown in Figure 1, encoder in current models is trained in two different ways: either jointly trained with decoder or solely trained through a reconstruction task. In the former end-to-end training scenario, it’s hard to evaluate encoder performance separately. So we mainly focus on the latter, in which case the encoder is trained through an encoder-decoder framework to reconstruct input cognitive signals. The decoder here doesn’t refer to the decoder for text generation. It’s similar to the structure of the encoder and will be abandoned once the encoder is trained. Although a proper evaluation index of the encoder’s representation ability is missing, validation loss is used to measure the influence of data contamination.
@@
We test different splitting methods on two cognitive datasets. The validation loss of encoder is shown in Figure 4. For fMRI, influenced by leakage of cognitive signals, the validation loss of method (b), (c), (d), (e) keeps dropping even with long training epochs. The encoder is actually overfitting and degrading. For method (a) and (f) without cognitive signal leakage, the validation loss quickly rises after reaching the lowest point with a few epochs, satisfying the basic rule of machine learning. For EEG, we find validation loss keeps dropping for all methods even with very long training epochs, regardless of cognitive signal leakage or not. We think the poor spatial resolution of EEG signal might lead to this phenomenon.
4.4.2 Decoder
@@
All state-of-the-art models choose pre-trained language model BART (Lewis et al., 2020) as decoder. On one hand, the powerful auto-regressive decoder is able to achieve fluent sentence-level open vocabulary text generation. On the other hand, if data contamination occurs, due to the feature of auto-regressive generation, the decoder will generate memorized text given the first few words, which is obviously an act of cheating.
@@
The influence of text stimuli leakage on decoder is detected through BLEU scores (Papineni et al., 2002) and ROUGE-1 scores (Lin, 2004), which measure text similarity between generated text and ground truth. If evaluation indicators keep improving as training epochs increase, we believe part of the test set is leaked into training set and the model is overfitting. For fMRI signal, we test five current dataset splitting methods under different training settings. As shown in Table 2, we test two kinds of UniCoRN models. One is UniCoRN with finely tuned hyper-parameters claimed in the original paper, and the other is $*{}^{*}$ with a randomly initialized encoder. Empirically, the former will perform much better than the latter. However, in method (a), (c), (d), due to text stimuli leakage, if we reduce the learning rate and extend training epochs, $*{}^{*}$ performs much better than UniCoRN and its performance keeps rising with longer training epochs. As to method (b) and (e) with no text stimuli leakage, changing training epochs or learning rates makes no obvious difference to model performance. For EEG signal, the conclusion is similar as shown in Table 3. For method (a) and (c) with text stimuli leakage, model performance keeps rising with longer training epochs. For method (d) without text stimuli leakage, both models reach optimal performance after the first few rounds of training epochs.
4.5 Fair Benchmark
@@
We evaluate current SOTA models for brain-to-text decoding under our dataset splitting method and release a fair benchmark. UniCoRN is tested for both fMRI and EEG decoding, EEG2Text model is tested for EEG decoding. The results are listed in Table 4. For EEG dataset, UniCoRN achieves higher results in BLEU-2,3,4 while EEG2Text is better in BLEU-1 and ROUGE-1.
@@ 5 Conclusion
In this paper, we explore a controversial topic: Due to the complexity of cognitive datasets, no consensus has been reached on how to split the dataset for training, validating, and testing in brain-to-text decoding. We analyze current dataset splitting methods and find serious data contamination largely exaggerates model performance and leads to poor generalization. Sufficient experiments and analysis are conducted to prove the existence of the above phenomena. Moreover, we propose a new dataset splitting method which can avoid both cognitive signal and text stimuli leakage to fix this problem. Current state-of-the-art models are reevaluated under this setting and a fair benchmark is released for further research in the domain.
@@
The "Narratives" dataset and the ZuCo dataset provide researchers with precise cognitive signal resources stimulated by text or voice. However, in brain-to-text decoding task, both subject’s cognitive signals and text stimuli in validation and test set need to be invisible to training set, which makes splitting these public datasets difficult. Our proposed dataset splitting method meets the above requirements at the expense of discarding some data in the dataset. We recommend future datasets in this domain follow these guidelines. The division of the training set, validation set, and test set should be provided when the dataset is released. Besides, we suggest hiring new subjects with unique stimuli for validation set and test set, which is good for testing the generalization ability of models without loss of data (Tang et al., 2023). What’s more, we find existing models rely more on the strong auto-regressive decoder to achieve good generation quality. The encoder is of limited use in all SOTA models, which might become a research point in the future.
@@
In this paper, we introduce a new dataset splitting method to avoid data contamination for decoding cognitive signals to text task. Experiments are conducted on public accessible cognitive datasets Narratives and ZuCo1.0 with the authorization from their respective maintainers. Both datasets have been de-identified by dataset providers and used for researches only.
Acknowledgements
This research is supported by the National Natural Science Foundation of China (No.62106105), the CCF-Baidu Open Fund (No.CCF-Baidu202307), the CCF-Zhipu AI Large Model Fund (No.CCF-Zhipu202315), the Scientific Research Starting Foundation of Nanjing University of Aeronautics and Astronautics (No.YQR21022), and the High Performance Computing Platform of Nanjing University of Aeronautics and Astronautics. We appreciate JD.com for providing powerful GPUs for experiments.
@@ Appendix A Implementation Details
More details in experiments are supplemented in this section. We perform the same filtering steps to Narratives dataset as UniCoRN paper (Xi et al., 2023) and the same filtering steps to ZuCo1.0 as EEG2Text paper (Wang and Ji, 2022). In CSLR and TSLR calculation, the number of four different seeds are set as $1,2,3,4$ respectively. In signal reconstruction task for encoder of UniCoRN, the batch size of EEG and fMRI data is 512 and 320 respectively. The learning rate is set as 1e-4 and 1e-3 separately as the author claimed in the original paper. In the fair benchmark, for fMRI data, encoder of UniCoRN is trained through 1e-4 learning rate and decaying to 1e-6 finally for 30 training epochs. Decoder is trained through 1e-4 learning rate and decaying to 1e-6 finally for 10 training epochs with 90 batch size. Sample length $L$ is set as 10 for all experiments related to fMRI. For EEG data, EEG2Text model is trained with 1e-6 learning rate for 80 epochs. UniCoRN model is trained with the same settings as fMRI data.
Appendix B Our Dataset Splitting Method
In this part, we release the pseudo-code of two dataset splitting methods for EEG and fMRI signal. As shown in Figure 3, our proposed dataset splitting method consists of four steps. The blue lines stand for the situation of original dataset. The main difference between two methods lies in the how $\mathcal{G}_{2}$ is generated. We always choose the side with fewer nodes in bipartite graph $\mathcal{G}_{1}$ to perform $\mathcal{G}_{2}$ generation. For example, in Algorithm 1 where we assert $|\mathcal{U}|<|\mathcal{V}|$, the adjacency matrix is initialized as $M\times N$. In Algorithm 2 where $|\mathcal{V}|<|\mathcal{U}|$, the adjacency matrix is initialized as $N\times K$. All hypotheses are based on analysis of cognitive datasets.

### v3

@@ 3.3 Cross-Subject Splitting Criterion
To eliminate data leakage from both brain signal leakage and text stimuli leakage, we split the dataset by $\left\langle S_{i},T_{j}\right\rangle$ pairs as shown in (f) of Figure 2. Since EEG and fMRI are different in the composition of dataset, we treat them separately and propose two data splitting methods. As to EEG dataset where $F_{ij}$ and $T_{j}$ form a sample, we consider a bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$ where $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{T_{j}\}_{j=1}^{M}$. $\mathcal{E}$ is the edge between node in $\mathcal{U}$ and node in $\mathcal{V}$, indicating $\left\langle S_{i},T_{j}\right\rangle$ pair in the dataset. $N$ is the total number of subjects and $M$ is the total number of unique text stimuli. We assert $M>N$, so $e=(u,v)\in\mathcal{E}$ exists for every $v\in\mathcal{V}$, as each text stimuli is listened or read by at least one subject. As shown in step 2 of Figure 3, first we pick one edge for each node $v\in\mathcal{V}$ and build a new bipartite graph $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$. Then following step 3, we split graph $\mathcal{G}_{2}$ by subject $\mathcal{U}$ with the given splitting ratio and form three disjoint graphs $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$. In step 4, some edges satisfying zero data leakage condition are not included in the graph. We add these edges to corresponding graphs, extending each graph $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ to its maximally scalable state and finishing the dataset splitting process.
@@
The sample in fMRI dataset is formatted in $F_{ij}=concat(f_{ij},f_{i,j+1},\ldots,f_{i,j+L-1})$ and $T_{j}=concat(s_{j},s_{j+1},\ldots,s_{j+L-1})$. If we follow the same process as EEG, text stimuli leakage will occur in the overlapping part of two samples, when one sample is assigned to training set and the other is assigned to validation or test set. We propose a simple solution that achieves the balance between abandoning as little data as possible and ensuring zero data leakage. Instead of $\left\langle S_{i},T_{j}\right\rangle$ pair, we consider $\left\langle S_{i},M_{k}\right\rangle$ pair and apply the above-mentioned algorithm. More details and pseudo-code are available in Appendix B.
@@
Table 2: Generation quality of UniCoRN model for fMRI under different training settings. Here UniCoRN indicates the encoder of UniCoRN is randomly initialized instead of pre-trained through signal reconstruction task.
4 Experimental Settings
We test some SOTA brain-to-text decoding models on two popular cognitive datasets Narratives (Nastase et al., 2021) and ZuCo (Hollenstein et al., 2018). Dataset details are introduced in Appendix A. Comprehensive experiments are conducted to evidence the existence of the following phenomena: (1) Brain signals and text stimuli in test set leak into training set in all current dataset splitting methods. (2) The model’s generalization ability, particularly that of the auto-regressive decoder, has been overestimated due to data leakage. Because the number of tasks in EEG dataset is too small and method (e) makes no difference to EEG as method (d), we only consider method (a), (c), (d) for EEG.
4.1 Implementation
We follow the same settings of UniCoRN (Xi et al., 2023) and EEG2Text (Wang and Ji, 2022), except all the datasets are split to the ratio of 8:1:1 for fair comparison. All experiments are conducted on NVIDIA A100-SXM4-40GB GPUs. More details are shown in Appendix A.
@@ 4.1 Implementation
$N=1$ $N=2$ $N=3$ $N=4$ $R$ $P$ $F$
@@
4.2 Data Leakage Metrics
We have analyzed two kinds of data leakage: brain signal leakage and text stimuli leakage in section 3. In this part, we will quantify two kinds of data leakage through experiments.
To better illustrate the extent of data leakage of different data splitting methods, we design two novel evaluation metrics named Brain Signal Leakage Rate (BSLR) and Text Stimuli Leakage Rate (TSLR) for detecting brain signal leakage and text stimuli leakage. Note that the situation for validation set is the same as test set, so we only consider test set in experiments. BSLR indicates the average percentage of each subject’s brain signals in test set appearing in training set, which could be formulated as
@@ 4.2 Data Leakage Metrics
where $N$ stands for the total number of subjects in test set. ||$\cdot$|| ⋅ | stands for the cardinality of a set. Function $\min(\cdot,\cdot)$ is applied to make sure for each subject the data leakage rate is less than one.
The definition of TSLR is somewhat different for EEG signal and fMRI signal. As to EEG signal where brain signals are sampled continuously, it’s easy to match certain text stimuli with corresponding signals. Its TSLR is similar to BSLR, which indicates the average percentage of certain text in test set appearing in training set. TSLR for EEG data can be calculated through
@@
The fMRI signal is sampled discretely with a deterministic interval TR, making it hard to acquire signals corresponding to text. Previous methods instead concatenated continuous fMRI frames of certain length with their corresponding text segments as training samples. As a result, we consider the average percentage of the same text segments in test set appearing in training set as TSLR for fMRI signal. It can be formulated as
@@
5 Results and Analysis
5.1 Verification for Data Leakage
We test current data splitting methods and our method on fMRI dataset “Narratives” and EEG dataset ZuCo. Considering the influence of randomness in splitting, we randomly select four seeds for experiments. The results are shown in Table 1, and are consistent with theoretical analysis. For fMRI, current methods apart from method (a) suffer from brain signal leakage, while method (a) has serious text stimuli leakage. Method (b) gets no text stimuli leakage but has slight brain signal leakage. The situation for EEG is similar to that of fMRI. Apart from our proposed method (f), there is no way to achieve zero brain signal leakage and text stimuli leakage at the same time.
5.2 Damage of Data Leakage
@@
Brain signal leakage and text stimuli leakage will damage brain-to-text decoding models from both the encoder side and decoder side.
Effect on Encoder
@@
The encoder of current brain-to-text decoding models can be trained in two ways: either jointly with the decoder or solely pre-trained through a reconstruction task. In the former end-to-end training scenario, it is hard to evaluate encoder performance separately. So we mainly focus on the latter, in which case the encoder is pre-trained through an encoder-decoder framework to reconstruct input brain signals. The decoder here does not refer to the decoder for text generation. It is similar to the structure of the encoder and will be abandoned once the encoder is pre-trained. Since a proper evaluation index of the encoder’s representation ability is missing, validation loss is applied to measure the effect of data leakage.
@@
We test different splitting methods on two cognitive datasets. The validation loss of encoder is shown in Figure 4. For fMRI, influenced by leakage of brain signals, the validation loss of method (b), (c), (d), (e) keeps dropping even with long training epochs. The encoder is actually overfitting and degrading. For method (a) and (f) without brain signal leakage, the validation loss quickly rises after reaching the lowest point with a few epochs, satisfying the basic rule of machine learning. For EEG, we find validation loss keeps dropping for all methods even with very long training epochs, regardless of brain signal leakage or not. We think the poor spatial resolution of EEG signal might lead to this phenomenon.
Effect on Decoder
@@
All SOTA models choose the pre-trained language model BART (Lewis et al., 2020) as decoder. The powerful model is able to achieve fluent open vocabulary text generation. However, if data leakage occurs, due to the auto-regressive generation that calculates the probability of current token based on all previous tokens, the decoder will generate seen text given the first few words, and fail to generalize to unseen text.
@@
The influence of text stimuli leakage on decoder is detected through BLEU scores (Papineni et al., 2002) and ROUGE-1 scores (Lin, 2004), which measure text similarity between generated text and ground truth. If evaluation indicators keep improving as training epochs increase, we believe part of the test set is leaked into training set and the model is overfitting. For fMRI signal, we test five current dataset splitting methods under different training settings. As shown in Table 2, we test two kinds of UniCoRN models. One is UniCoRN with finely tuned hyper-parameters claimed in the original paper, and the other is UniCoRN with a randomly initialized encoder. Empirically, the former will perform much better than the latter. However, in method (a), (c), (d), due to text stimuli leakage, if we reduce the learning rate and extend training epochs, UniCoRN performs much better than UniCoRN and its performance keeps rising with longer training epochs. As to method (b) and (e) with no text stimuli leakage, changing training epochs or learning rates makes no obvious difference to model performance. For EEG signal, the conclusion is similar as shown in Table 3. For method (a) and (c) with text stimuli leakage, model performance keeps rising with longer training epochs. For method (d) without text stimuli leakage, both models reach optimal performance after the first few rounds of training epochs.
5.3 A Fair Benchmark
@@
We re-evaluate current SOTA models for brain-to-text decoding under our cross-subject data splitting method and release a fair benchmark. UniCoRN is tested for both fMRI and EEG decoding, EEG2Text model is tested for EEG decoding. The results are listed in Table 4. For EEG dataset, UniCoRN achieves higher results in BLEU-2,3,4 while EEG2Text is better in BLEU-1 and ROUGE-1.
@@ 6 Conclusion
In this paper, we conduct a comprehensive study on existing cross-subject data splitting methods, and evidence that all these methods suffer from data leakage problem. Such data leakage largely exaggerates model performance and leads to poor generalization. To fix this issue, we propose a cross-subject data splitting criterion for brain-to-text decoding, aiming to improve the utilization efficiency of cognitive dataset and the generalization ability of decoding models. Experiments are conducted on fMRI and EEG dataset respectively. Current SOTA models are re-evaluated under this proposed splitting method and a fair benchmark is released for further research in this domain.
@@
The “Narratives” dataset and the ZuCo dataset provide researchers with precise brain signal resources stimulated by text or voice. However, in the brain-to-text decoding task, both subject’s brain signals and text stimuli in the validation and test set need to be invisible to the training set, which makes splitting these public datasets difficult. Our proposed dataset splitting method meets the above requirements at the expense of discarding some data in the dataset. We recommend future datasets in this domain follow these guidelines. The division of the training set, validation set, and test set should be provided when the dataset is released. Besides, we suggest hiring new subjects with unique stimuli for the validation set and test set, which is good for testing the generalization ability of models without loss of data. What’s more, we find existing models rely more on a strong auto-regressive decoder to achieve good generation quality. The encoder is of limited use in all SOTA models. And we also notice in experiments that the encoder of EEG2Text keeps overfitting whether with or without brain signal leakage. We leave it to research in the future.
@@
In this paper, we introduce a new dataset splitting method to avoid data leakage for decoding brain signals to text task. Experiments are conducted on the publicly accessible cognitive datasets “Narratives” and ZuCo1.0 with the authorization from their respective maintainers. Both datasets have been de-identified by dataset providers and used for researches only.
@@ Appendix A Implementation Details
[…]
More details in experiments are supplemented in this section. We perform the same filtering steps to “Narratives” dataset as UniCoRN paper (Xi et al., 2023) and the same filtering steps to ZuCo1.0 as EEG2Text paper (Wang and Ji, 2022). In BSLR and TSLR calculation, the number of four different seeds are set as $1,2,3,4$ respectively. In signal reconstruction task for encoder of UniCoRN, the batch size of EEG and fMRI data is 512 and 320 respectively. The learning rate is set as 1e-4 and 1e-3 separately as the author claimed in the original paper. In the fair benchmark, for fMRI data, encoder of UniCoRN is trained through 1e-4 learning rate and decaying to 1e-6 finally for 30 training epochs. Decoder is trained through 1e-4 learning rate and decaying to 1e-6 finally for 10 training epochs with 90 batch size. Sample length $L$ is set as 10 for all experiments related to fMRI. For EEG data, EEG2Text model is trained with 1e-6 learning rate for 80 epochs. UniCoRN model is trained with the same settings as fMRI data.
Appendix B Details of Cross-Subject Splitting
For fMRI dataset, we consider a bipartite graph $\mathcal{G}_{1}=(\mathcal{U},\mathcal{V},\mathcal{E})$ where $\mathcal{U}=\{S_{i}\}_{i=1}^{N}$, $\mathcal{V}=\{M_{k}\}_{k=1}^{M}$. $\mathcal{E}$ is the edge between node in $\mathcal{U}$ and node in $\mathcal{V}$, indicating $\left\langle S_{i},M_{k}\right\rangle$ pair in the dataset. $N$ is the total number of subjects and $M$ is the total number of tasks. We assert $M<N$, so $e=(u,v)\in\mathcal{E}$ exists for every $v\in\mathcal{V}$, as each text stimuli is listened or read by at least one subject. As shown in step 2 of Figure 3, first we pick one edge for each node $v\in\mathcal{V}$ and build a new bipartite graph $\mathcal{G}_{2}=(\mathcal{U},\mathcal{V},\mathcal{E}^{\prime})$. Then following step 3, we split graph $\mathcal{G}_{2}$ by subject $\mathcal{U}$ with the given splitting ratio and form three disjoint graphs $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$. In step 4, some edges satisfying zero data leakage condition are not included in the graph. We add these edges to corresponding graphs, extending each graph $\mathcal{G}_{train},\mathcal{G}_{val},\mathcal{G}_{test}$ to its maximally scalable state and finishing the dataset splitting process.
We also release the pseudo-code of two dataset splitting methods for EEG and fMRI signal. As shown in Figure 3, our proposed dataset splitting method consists of four steps. The blue lines stand for the situation of original dataset. The main difference between two methods lies in the how $\mathcal{G}_{2}$ is generated. We always choose the side with fewer nodes in bipartite graph $\mathcal{G}_{1}$ to perform $\mathcal{G}_{2}$ generation. For example, in Algorithm 1 where we assert $|\mathcal{U}|<|\mathcal{V}|$, the adjacency matrix is initialized as $M\times N$. In Algorithm 2 where $|\mathcal{V}|<|\mathcal{U}|$, the adjacency matrix is initialized as $N\times K$. All assertions are based on analysis of cognitive datasets.